Let S be the set of all real numbers except −1. Define * on S by


a * b = a + b + ab


Show that〈 S,*〉is a group and solve 2 * x * 3 = 7 in S



.



1
Expert's answer
2021-02-19T15:21:20-0500

Given that "a \neq-1, b \neq -1 \in \mathbb{R}" , we know that "(\mathbb{R},+)" and "(\mathbb{R},.)" are groups. Hence "S" is closed.

Next, we show that S is associative

Let "a,b, c \in S"

Consider "(a*b)*c = (a+b+ab)*c = a+b+ab + c + (a+b+ab)c"

"= a+b+c +ab +ac + bc +abc= a+ab+ac+abc +b+c+bc = a + a(b+c+bc) + (b+c+bc)"

"=a+ a(b*c) + (b*c) = a + (b*c) + a(b*c) = a*(b*c)"

Next we show that S has an identity element. Inorder to do this, we suppose there is an "e \in S ~ \ni a*e =a" i.e "a + e +ae = a \\
\implies e +ae = 0\\ \implies e(1+a)=0"

Either "e =0" or "1+a = 0"

If "1+a =0" ,then "a \notin S \therefore e= 0"

(Since "(\mathcal{R},+)" and "(\mathcal{R}, .)" are Abelian groups, we have that

"a*e=a=e*a)"

Next we show that every non zero element of "S" has an inverse.

Let "a,b \in S"

Consider,

"a+b+ab =e=0\\
b+ab = -a\\
b(1+a) = -a\\
b = \frac{-a}{1+a}"

By definition of elements of "S," we have that "a\neq -1" . Hence for every non-zero "a\in S" , "a^{-1} = \frac{-a}{1+a}"

Hence "\langle S, * \rangle" is a group

"(2*x)*3 = 7\\ (2+x+2x)*3 =7\\ (2+3x)*3=7\\ 2+3x +3+(2+3x)3 = 7\\ 5+ 3x + 6 + 9x =7\\12x = 7-11\\ 12x = -4 \\ x = \frac{-4}{12}\\ x = \frac{-1}{3}"